write the standard form of an ellipse.
x^2+4y^2+6x-91=0

Answer by lwsshak3(11628)

write the standard form of an ellipse:
x^2+4y^2+6x-91=0
x^2+6x+4y^2=91
complete the square:
(x^2+6x+9)4y^2=9+91
(x+3)^2+4y^2=100
(x+3)^2/100+y^2/25=1
This is and ellipse with horizontal major axis
Its standard form:(x-h)^2/a^2+(y-k)^2/b^2 a>b,(h,k)=(x,y) coordinates of the center: